OpenStudy (anonymous):

Please help people... Simplify 5^2x - 24 * 5^x = 25

OpenStudy (anonymous):

i assume "simplify" actually means "solve for \(x\) in this case, right?

OpenStudy (anonymous):

yessh :)

OpenStudy (anonymous):

start with \[5^{2x}-24\times 5^x-25=0\] factor as \[(5^x-25)(5^x+1)=0\]

OpenStudy (anonymous):

if that is confusing, note that \(5^{2x}=(5^x)^2\) and think of it as \[u^2-24u-25=0\] where \(u=5^x\) then it will be easier to see how to factor it

OpenStudy (anonymous):

uhh yah but cudnt yer just multiply the 5^2x and da 5^x and get 5^3x?

OpenStudy (anonymous):

once you have it factored as \[u^2-24y=u-25=(u-25)(u+1)\] then replace \(u\) by \(5^x\) to get \[(5^x-25)(5^x+1)=0\] and solve for \(x\) notice that \(5^x+1=0\) has not solutions, because \(5^x>0\) always

OpenStudy (anonymous):

it is not a multiplication, so no you cannot

OpenStudy (anonymous):

but they arent any brackets, so?

OpenStudy (anonymous):

coz im guessin if they wer brackets like (5^2x -24) * 5^x then u cudnt multiply..